how would you solve this 4x4 system of equations:
4w+ x+2y-3z=-16
-3w+3x- y+4z= 20
-w+2x+5y+ z= -4
5w+4x+3y- z=-10

Answer by Edwin McCravy(20086)

(1)      4w +  x + 2y - 3z = -16
(2)     -3w + 3x -  y + 4z =  20
(3)      -w + 2x + 5y +  z =  -4
(4)      5w + 4x + 3y -  z = -10

Pick a letter to eliminate.  I pick z.

Add equations (3) and (4) as they are

(3)      -w + 2x + 5y +  z =  -4
(4)      5w + 4x + 3y -  z = -10
--------------------------------
         4w + 6x + 8y      = -14

That can be simplified by dividing through by 2

(5)     2w + 3x + 4y      =  -7

Multiply (4) by 4 [to add to (2) to eliminate z]

        20w + 16x + 12y - 4z = -40
(2)     -3w +  3x -   y + 4z =  20
----------------------------------
(6)     17w + 19x + 11y      = -20       

Multiply (3) by 3 [to add to (1) to eliminate z]

        -3w + 6x + 15y + 3z =  -12
(1)      4w +  x +  2y - 3z =  -16
----------------------------------
(7)       w + 7x + 17y      =  -28

Now we have reduced the system to 3 equations in 3 unknowns:

(5)      2w +  3x +  4y  =  -7
(6)     17w + 19x + 11y  = -20
(7)       w +  7x + 17y  = -28

Pick a letter to eliminate.  I pick w.

Multiply (7) by -2 [to add to (5) to eliminate w]

        -2w - 14x - 34y  =  56
(5)      2w +  3x +  4y  =  -7
------------------------------
(8)           -11x - 30y  =  49

Multiply (7) by -17 [to add to (6) to eliminate w]
     
       -17w - 119x - 289y  = 476
(6)     17w +  19x +  11y  = -20
--------------------------------
             -100x - 278y  = 456

That can be simplified by dividing through by 2
            
(9)           -50x - 139y  = 228

Now we have reduced the system to 2 equations in 2 unknowns:

(8)           -11x -  30y  =  49
(9)           -50x - 139y  = 228

Pick a letter to eliminate.  I pick x.

the least common multiple of 11 and 50 is 550. So we multiply
(8) by 50, and(9) by -11 so that the x's will cancel:

            -550x - 1500y  =  2450
             550x + 1529y  = -2508
           -------------------------
                      29y  =   -58
(10)                    y = -2

Substitute in (8)

           -11x -  30(-2)  =  49
                 -11x + 60 =  49
                      -11x = -11
(11)                     x = 1

Substitute x = 1 and y = -2 in (5)

           2w +  3x +  4y  = -7
         2w + 3(1) + 4(-2) = -7
                2w + 3 - 8 = -7
                    2w - 5 = -7
                        2w = -2
                         w = -1

Substitute x = 1, y = -2 and w = -1 in (3)
 
 -(-1) + 2(1) + 5(-2) +  z = -4
            1 + 2 - 10 + z = -4
                    -7 + z = -4
                         z = 3

(w,x,y,z) = (-1,1,-2,3)

Edwin